# Closure Discovery for Coarse-Grained Partial Differential Equations using Multi-Agent Reinforcement Learning

### v1

@@ 3.7 Computational Complexity
We note that the computational complexity of the CGS w.r.t. the number of discretization points scales with $\mathcal{O}(\tilde{N}_{x}\tilde{N}_{y})$. As one forward pass through the FCN also scales with $\mathcal{O}(\tilde{N}_{x}\tilde{N}_{y})$ the same is true for CNN-MARL. The FGS employs a finer grid, which leads to a computational cost that scales with $\mathcal{O}(d_{t}d^{2}\tilde{N}_{x}\tilde{N}_{y})$. This indicates a scaling advantage for CNN-MARL compared to a FGS. Based on these considerations and as shown in the next section in the experiments, CNN-MARL is able to compress some of the computations that are performed on the fine grid as it is able to significantly improve the CGS while keeping the execution time below that of the FGS.
@@ 4.1 Advection Equation
First we apply CNN-MARL to the 2D advection equation [^3] The code to reproduce all results in this paper can be found at (URL added upon publication).:
@@
For the FGS, this domain is discretized using $N_{x}=N_{y}=256$ discretization points in each dimension. To guarantee stability, we employ a time step that ensures that the Courant–Friedrichs–Lewy (CFL) (Courant et al., 1928) condition is fulfilled. The spatial derivatives are calculated using central differences and the time stepping uses the fourth-order Runge-Kutta scheme (Quarteroni & Valli, 2008). The FGS is fourth order accurate in time and second order accurate in space.
We construct the CGS by employing the subsampling factors $d=4$ and $d_{t}=4$. Spatial derivatives in the CGS use an upwind scheme and time stepping is performed with the forward Euler method, resulting in first order accuracy in both space and time. All simulation parameters for CGS and FGS are summarized in Table 3.
@@ 4.1.1 Initial Conditions
In order to prevent overfitting and promote generalization, we design the initializations of $\psi,u$ and $v$ to be different for each episode while still fulfilling the PBCs and guaranteeing the incompressibility of the velocity field. The velocity fields are sampled from a distribution $\mathcal{D}^{Vortex}_{Train}$ by taking a linear combination of Taylor-Greene vortices and an additional random translational field. Further details are provided in Section C.2.1. For visualization purposes, the concentration of a new episode is set to a random sample from the MNIST dataset (Deng, 2012) that is scaled to values in the range $[0,1]$. In order to increase the diversity of the initializations, we augment the data by performing random rotations of ±$\pm 90^{\circ}$ in the image loading pipeline.
Figure 3: Results of CNN-MARL applied to the advection equation. The relative MAEs are computed over 100 simulations with respect to (w.r.t.) the FGS. The shaded regions correspond to the respective standard deviations. The violin plot on the right shows the number of simulation steps until the relative MAE w.r.t. the FGS reaches the threshold of $1\%$.
@@ 4.1.2 Training
We train the framework for a total of 2000 epochs and collect 1000 transitions in each epoch. More details regarding the training process as well as the training time are provided in Appendix C.
We note that the amount and length of episodes varies during the training process: The episodes are truncated based on the relative Mean Absolute Error (MAE) defined as $\text{MAE}(\psi^{n},\tilde{\psi^{n}})=\frac{1}{\tilde{N}_{x}\cdot\tilde{N}_{y}}\sum_{i,j}|\psi^{n}_{ij}-\tilde{\psi}^{n}_{ij}|/\psi_{max}$ between the CGS and FGS concentrations. Here, $\psi_{max}$ is the maximum observable value of the concentration $\psi$. If this error exceeds the threshold of $1.5\%$, the episode is truncated. This ensures that during training, the CGS and FGS stay close to each other so that the reward signal is meaningful. As the agents become better during the training process, the mean episode length increases as the two simulations stay closer to each other for longer.
We designed this adaptive training procedure in order to obtain stable simulations.
@@ 4.1.4 In- and Out-of-Distribution MAE
Figure 4: Example run comparing CGS, CNN-MARL and FGS with the same ICs. The concentration is sampled from the test set and the velocity components are randomly sampled from $\mathcal{D}^{Vortex}_{Train}$. Note that the agents of the CNN-MARL method have only been trained up to $n=100$. However, they qualitatively improve the CGS simulation past that point.
We develop metrics for CNN-MARL by running 100 simulations of 50 time steps each with different ICs. For the in-distribution case, the concentrations are sampled from the MNIST test set and the velocity fields are sampled from $\mathcal{D}^{Vortex}_{Train}$. To quantify the performance on out-of-distribution ICs, we also run evaluations on simulations using the Fashion-MNIST dataset (F-MNIST) (Xiao et al., 2017) and a new distribution, $\mathcal{D}^{Vortex}_{Test}$, for the velocity fields. The latter is defined in Section C.2.2. The resulting error metrics of CGS, ACGS and CNN-MARL w.r.t. the FGS are collected in Table 1. A qualitative example of the CGS, FGS and the CNN-MARL method after training is presented in Figure 4. The example shows that CNN-MARL is able to compensate for the dissipation that is introduced by the first order scheme and coarse grid in the CGS.
@@
CNN-MARL $\textbf{1.46}\pm 0.33$ $\textbf{2.12}\pm 0.57$ $\textbf{1.58}\pm 0.37$ $\textbf{2.04}\pm 0.56$
@@
CNN-MARL -53% -34% -58% -36%
CNN-MARL reduces the relative MAE of the CGS after 50 steps by $30\%$ or more in both in- and out-of-distribution cases. This shows that the agents have learned a meaningful correction for the truncation errors of the numerical schemes in the coarser grid.
CNN-MARL also outperforms the ACGS w.r.t. the MAE metric which indicates that the learned corrections emulate a higher-order scheme. This indicates that the proposed methodology is able to emulate the unresolved dynamics and is a suitable option for complementing existing numerical schemes.
We note the strong performance of our framework w.r.t. the out-of-distribution examples. For both unseen and out-of-distribution ICs as well as parameters, the framework was able to outperform CGS and ACGS. In our opinion, this indicates that we have discovered an actual model of the forcing terms that goes beyond the training scenarios.
@@ 4.1.5 Evolution of Numerical Error
The plots of the evolution of the relative error show that CNN-MARL is able to improve the CGS for the entire range of a 400-step rollout, although it has only been trained for 100 steps. This implies that the agents are seeing distributions of the concentration that have never been encountered during training and are able to generalize to these scenarios. When measuring the duration of simulations for which the relative error stays below $1\%$, we observe that the CNN-MARL method outperforms both ACGS and CGS, indicating that the method is able to produce simulations with higher long term stability than CGS and ACGS. We attribute this to our adaptive scheme for episode truncation during training as introduced in Section 4.1.2 and note that the increased stability can be observed well beyond the training regime.
@@ 4.1.6 Interpretation of Actions
We are able to derive the optimal update rule for the advection CGS that negates the errors introduced by the numerical scheme (see Appendix D for the derivation):
$\small{\tilde{\epsilon}^{n}=\frac{\widetilde{\Delta t}}{2}\tilde{\psi}_{tt}^{n}+|\tilde{u}^{n}|\frac{\widetilde{\Delta x}}{2}\tilde{\psi}_{xx}^{n}+|\tilde{v}^{n}|\frac{\widetilde{\Delta y}}{2}\tilde{\psi}_{yy}^{n}+\mathcal{O}(\widetilde{\Delta t}^{2},\widetilde{\Delta x}^{2},\widetilde{\Delta y}^{2})}$ (15)
In order to further quantify the similarity between the numerical estimates of $-\widetilde{\Delta t}\tilde{\epsilon}^{n-1}_{i,j}$ and the taken action, we compute the Pearson product-moment correlation coefficient for 100 samples. The results are presented in Table 2 and show that the learned actions as well as the optimal action of the CGS are highly correlated for all different combinations of seen and unseen ICs and parameters.
MNIST F-MNIST
$\mathcal{D}^{Vortex}_{Train}$ $0.82\pm 0.05$ $0.70\pm 0.08$
$\mathcal{D}^{Vortex}_{Test}$ $0.82\pm 0.03$ $0.72\pm 0.08$
Figure 6: Results of CNN-MARL applied to the Burgers’ equation. The relative MAEs are computed over 100 simulations with respect to the FGS. The shaded regions correspond to the respective standard deviations. The violin plot on the right shows the number of simulation steps until the relative MAE w.r.t. the FGS reaches the threshold of $10\%$.
Additionally, we note that the truncation error contains a second-order temporal derivative. At first glance, it might seem surprising that the model would be able to predict this temporal derivative as the agents can only observe the current time step. However, the observation contains both the PDE solution, i.e. the concentration for the present examples, as well as the parameters of the PDE, i.e. the velocity fields. Thus, both the concentration and velocity field are passed into the FCN and due to the velocity fields enough information is present to infer this temporal derivative.
@@ 4.2 Burgers’ Equation
Here, $\boldsymbol{\psi}:=(u,v)$ consists of both velocity components and the PDE has the single input parameter $C=\nu$. As for the advection equation, we assume periodic boundary conditions on the domain $\Omega=[0,1]\times[0,1]$. In comparison to the advection example, we are now dealing with two solution variables and thus $k=2$.
For the FGS, the aforementioned domain is discretized using $N_{x}=N_{y}=150$ discretization points in each dimension. Moreover, we again choose $\Delta t$ to fulfill the CFL condition for stability (see Table 4). The spatial derivatives are calculated using the upwind scheme and the forward Euler method is used for the time stepping (Quarteroni & Valli, 2008).
We construct the CGS by employing the subsampling factors $d=5$ and $d_{t}=10$. For the Burgers experiment, we apply a mean filter $K$ with kernel size $d\times d$ before the actual subsampling operation. The mean filter is used to eliminate higher frequencies in the fine grid state variables, which would lead to accumulating high errors. The CGS employs the same numerical schemes as the FGS here. This leads to first order accuracy in both space and time. All the simulation parameters for CGS and FGS are collected in Table 4.
In this example, where FGS and CGS are using the same numerical scheme, the CNN-MARL framework has to focus solely on negating the effects of the coarser discretization.
@@
For training and evaluation, we generate random, incompressible velocity fields as ICs (see Section C.2.1 for details) and set the viscosity $\nu$ to $0.003$. We observed that the training of the predictions actions can be improved by multiplying the predicted forcing terms with $\widetilde{\Delta t}$. This is consistent with our previous analysis in Section 4.1.6 as the optimal action is also multiplied but this factor. Again, we train the model for 2000 epochs with 1000 transitions each. The maximum episode length during training is set to $200$ steps and, again, we truncate the episodes adaptively, when the relative error exceeds $20\%$. Further details on training CNN-MARL for the Burgers’ equation can be found in Appendix C. Visualization of the results can be found in Section B.2 and the resulting relative errors w.r.t. the FGS are shown in Figure 6. The CNN-MARL method again improves the CGS significantly, also past the point of the 200 steps seen during training. Specifically, in the range of step 00 to step $100$ in which the velocity field changes fastest, we see a significant error reductions up to $-80\%$. When analyzing the duration for which the episodes stay under the relative error of $10\%$, we observe that the mean number of steps is improved by two order of magnitude, indicating that the method is able to improve the long term accuracy of the CGS.
5 Conclusion
We propose a novel methodology (CNN-MARL), for the automated discovery of closure models of coarse grained discretizations of time-dependent PDEs. We show that CNN-MARL develops a policy that compensates for numerical errors in a CGS of the 2D advection and Burgers’ equation. More importantly we find that, after training, the learned closure model can be used for predictions in extrapolative test cases. In turn, the agent actions are interpretable and highly correlated with the error terms of the numerical scheme.
@@ 5 Conclusion
CNN-MARL utilizes a multi-agent formulation with a FCN for both the policy and the value network. This enables the incorporation of both local and global rewards without necessitating individual neural networks for each agent. This setup allows to efficiently train a large number of agents without the need for decoupled subproblems. Moreover, the framework trains on rollouts without needing to backpropagate through the numerical solver itself.
A current drawback is the requirement of a regular grid. The framework would still be applicable to an irregular grid, but information about the geometry would need to be added to the observations. Moreover, for very large systems of interest the numerical schemes used are often multi-grid and the RL framework should reflect this. For such cases, we suggest defining separate rewards for each of the grids employed by the numerical solver.
We believe that the introduced framework holds significant promise for advancing the discovery of closure models in a wide range of systems governed by Partial Differential Equations as well as other computational methods such as coarse grained molecular dynamics.
Broader Impact Statement
In this paper, we have presented a MARL framework for systematic identification of closure models in coarse-grained PDEs. The anticipated impact pertains primarily to the scientific community, whereas the use in real-life applications would require additional adjustments and expert input.
Appendix A Adapted PPO Algorithm

### v3

@@ 3.6 Computational Complexity
We note that the computational complexity of the CGS w.r.t. the number of discretization points scales with $\mathcal{O}(\tilde{N}_{x}\tilde{N}_{y})$. As one forward pass through the FCN also scales with $\mathcal{O}(\tilde{N}_{x}\tilde{N}_{y})$ the same is true for Closure-RL. The FGS employs a finer grid, which leads to a computational cost that scales with $\mathcal{O}(d_{t}d^{2}\tilde{N}_{x}\tilde{N}_{y})$. This indicates a scaling advantage for Closure-RL compared to a FGS. Based on these considerations and as shown in the experiments, Closure-RL is able to compress some of the computations that are performed on the fine grid as it is able to significantly improve the CGS while keeping the execution time below that of the FGS.
@@ 4.1 Advection Equation
First we apply Closure-RL to the 2D advection equation [^3] The code to reproduce all results in this paper can be found at (URL added upon publication).:
@@
For the FGS, this domain is discretized using $N_{x}=N_{y}=256$ discretization points in each dimension. To guarantee stability, we employ a time step that ensures that the Courant–Friedrichs–Lewy (CFL) [43] condition is fulfilled. The spatial derivatives are calculated using central differences and the time stepping uses the fourth-order Runge-Kutta scheme [44]. The FGS is fourth order accurate in time and second order accurate in space.
We construct the CGS by employing the subsampling factors $d=4$ and $d_{t}=4$. Spatial derivatives in the CGS use an upwind scheme and time stepping is performed with the forward Euler method, resulting in first order accuracy in both space and time. All settings of numerical values used for the CGS and FGS are summarized in Table 2.
@@ 4.1.1 Initial Conditions
In order to prevent overfitting and promote generalization, we design the initializations of $\psi,u$ and $v$ to be different for each episode while still fulfilling the PBCs and guaranteeing the incompressibility of the velocity field. The velocity fields are sampled from a distribution $\mathcal{D}^{Vortex}_{Train}$ by taking a linear combination of Taylor-Greene vortices and an additional random translational field. Further details are provided in Section D.2.1. For visualization purposes, the concentration of a new episode is set to a random sample from the MNIST dataset [45] that is scaled to values in the range $[0,1]$. In order to increase the diversity of the initializations, we augment the data by performing random rotations of ±$\pm 90^{\circ}$ in the image loading pipeline.
Figure 2: Results of Closure-RL applied to the advection equation. The relative MAEs are computed over 100 simulations with respect to (w.r.t.) the FGS. The shaded regions correspond to the respective standard deviations. The violin plot on the right shows the number of simulation steps until the relative MAE w.r.t. the FGS reaches the threshold of $1\%$.
@@ 4.1.2 Training
We train the framework for a total of 2000 epochs and collect 1000 transitions in each epoch. More details regarding the training process as well as the training time are provided in Appendix D.
We note that the amount and length of episodes varies during the training process: The episodes are truncated based on the relative Mean Absolute Error (MAE) defined as $\text{MAE}(\psi^{n},\tilde{\psi^{n}})=\frac{1}{\tilde{N}_{x}\cdot\tilde{N}_{y}}\sum_{i,j}|\psi^{n}_{ij}-\tilde{\psi}^{n}_{ij}|/\psi_{max}$ between the CGS and FGS concentrations. Here, $\psi_{max}$ is the maximum observable value of the concentration $\psi$. If this error exceeds the threshold of $1.5\%$, the episode is truncated. This ensures that during training, the CGS and FGS stay close to each other so that the reward signal is meaningful. As the agents become better during the training process, the mean episode length increases as the two simulations stay closer to each other for longer. We designed this adaptive training procedure in order to obtain stable simulations.
We note that inspired by [24] we experimented with adding an additional domain-specific global term to the reward (see Appendix E for details). We did not see any improvements compared to our Closure-RL formulation which indicates that the proposed reward structure is sufficient to learn an accurate closure model.
@@ 4.1.4 In- and Out-of-Distribution MAE
Figure 3: Example run comparing CGS, Closure-RL and FGS with the same ICs. The concentration is sampled from the test set and the velocity components are randomly sampled from $\mathcal{D}^{Vortex}_{Train}$. Note that the agents of the Closure-RLmethod have only been trained up to $n=100$. However, they qualitatively improve the CGS simulation past that point.
We develop metrics for Closure-RL by running 100 simulations of 50 time steps each with different ICs. For the in-distribution case, the concentrations are sampled from the MNIST test set and the velocity fields are sampled from $\mathcal{D}^{Vortex}_{Train}$. To quantify the performance on out-of-distribution ICs, we also run evaluations on simulations using the Fashion-MNIST dataset (F-MNIST) [46] and a new distribution, $\mathcal{D}^{Vortex}_{Test}$, for the velocity fields. The latter is defined in Section D.2.2. The resulting error metrics of CGS, ACGS and Closure-RL w.r.t. the FGS are collected in Table 1. A qualitative example of the CGS, FGS and the Closure-RL method after training is presented in Figure 3. The example shows that Closure-RL is able to compensate for the dissipation that is introduced by the first order scheme and coarse grid in the CGS.
@@
Closure-RL $\textbf{1.46}\pm 0.33$ $\textbf{2.12}\pm 0.57$ $\textbf{1.58}\pm 0.37$ $\textbf{2.04}\pm 0.56$
@@
Closure-RL -53% -34% -58% -36%
Closure-RL reduces the relative MAE of the CGS after 50 steps by $30\%$ or more in both in- and out-of-distribution cases. This shows that the agents have learned a meaningful correction for the truncation errors of the numerical schemes in the coarser grid. Closure-RL also outperforms the ACGS w.r.t. the MAE metric which indicates that the learned corrections emulate a higher-order scheme. This indicates that the proposed methodology is able to emulate the unresolved dynamics and is a suitable option for complementing existing numerical schemes.
We note the strong performance of our framework w.r.t. the out-of-distribution examples. For both unseen and out-of-distribution ICs as well as PDEPs, the framework was able to outperform CGS and ACGS. In our opinion, this indicates that we have discovered an actual model of the forcing terms that goes beyond the training scenarios.
@@ 4.1.5 Evolution of Numerical Error
The plots of the evolution of the relative error show that Closure-RL is able to improve the CGS for the entire range of a 400-step rollout, although it has only been trained for 100 steps. This implies that the agents are seeing distributions of the concentration that have never been encountered during training and are able to generalize to these scenarios. When measuring the duration of simulations for which the relative error stays below $1\%$, we observe that the Closure-RL method outperforms both ACGS and CGS, indicating that the method is able to produce simulations with higher long term stability than CGS and ACGS. We attribute this to our adaptive scheme for episode truncation during training as introduced in Section 4.1.2 and note that the increased stability can be observed well beyond the training regime.
@@ 4.1.6 Interpretation of Actions
The intention of our closure scheme is to negate the errors introduced by the numerical scheme. For the advection CGS, we are able to show that the actions taken are indeed highly correlated with the optimal action based on the errors of the numerical scheme used. In Appendix G, we are reporting the derivation of this optimal action as well as the correlation with the actions actually taken by Closure-RL. The obtained correlation coefficients are between $0.7-0.82$ depending on the task.
@@ 4.2 Burgers’ Equation
Figure 4: Results of Closure-RL applied to the Burgers’ equation. The relative MAEs are computed over 100 simulations with respect to the FGS. The shaded regions correspond to the respective standard deviations. The violin plot on the right shows the number of simulation steps until the relative MAE w.r.t. the FGS reaches the threshold of $10\%$.
@@
Here, $\boldsymbol{\psi}:=(u,v)$ consists of both velocity components and the PDE has the single input PDEP $C=\nu$. As for the advection equation, we assume periodic boundary conditions on the domain $\Omega=[0,1]\times[0,1]$. In comparison to the advection example, we are now dealing with two solution variables and thus $k=2$.
@@
For the FGS, the aforementioned domain is discretized using $N_{x}=N_{y}=150$ discretization points in each dimension. Moreover, we again choose $\Delta t$ to fulfill the CFL condition for stability (see Table 3). The spatial derivatives are calculated using the upwind scheme and the forward Euler method is used for the time stepping [44]. We construct the CGS by employing the subsampling factors $d=5$ and $d_{t}=10$. For the Burgers experiment, we apply a mean filter $K$ with kernel size $d\times d$ before the actual subsampling operation. The mean filter is used to eliminate higher frequencies in the fine grid state variables, which would lead to accumulating high errors. The CGS employs the same numerical schemes as the FGS here. This leads to first order accuracy in both space and time. All numerical settings used for the CGS and FGS are collected in Table 3. In this example, where FGS and CGS are using the same numerical scheme, the Closure-RL framework has to focus solely on negating the effects of the coarser discretization. For training and evaluation, we generate random, incompressible velocity fields as ICs (see Section D.2.1 for details) and set the viscosity $\nu$ to $0.003$. We observed that the training of the predictions actions can be improved by multiplying the predicted forcing terms with $\widetilde{\Delta t}$. This is consistent with our previous analysis in Appendix G as the optimal action is also multiplied but this factor. Again, we train the model for 2000 epochs with 1000 transitions each. The maximum episode length during training is set to $200$ steps and, again, we truncate the episodes adaptively, when the relative error exceeds $20\%$. Further details on training Closure-RL for the Burgers’ equation can be found in Appendix D. Visualization of the results can be found in Section C.2 and the resulting relative errors w.r.t. the FGS are shown in Figure 4. The Closure-RL method again improves the CGS significantly, also past the point of the 200 steps seen during training. Specifically, in the range of step 00 to step $100$ in which the velocity field changes fastest, we see a significant error reductions up to $-80\%$. When analyzing the duration for which the episodes stay under the relative error of $10\%$, we observe that the mean number of steps is improved by two order of magnitude, indicating that the method is able to improve the long term accuracy of the CGS.
5 Conclusions
@@
We propose Closure-RL, a novel framework for the automated discovery of closure models of coarse grained discretizations of time-dependent PDEs. It utilizes a grid-based RL formulation with a FCN for both the policy and the value network. This enables the incorporation of local rewards without necessitating individual neural networks for each agent and allows to efficiently train a large number of agents. Moreover, the framework trains on rollouts without needing to backpropagate through the numerical solver itself. We show that Closure-RL develops a policy that compensates for numerical errors in a CGS of both the 2D advection and Burgers’ equation. More importantly we find that the learned closure model can be used for predictions in extrapolative test cases.
Further work may focus on extending the formulation to irregular grids by using a Graph Convolutional Network [47] instead of the FCN. Similarly to the current approach, we can place one agent at each discretization point whose receptive field now depends on the connection between the graph nodes. Moreover, for very large systems of interest the numerical schemes used are often multi-grid and the grid-based RL framework should reflect this. For such cases, we suggest defining separate rewards for each of the grids employed by the numerical solver. Additionally, Closure-RL may serve as a stepping-stone towards incorporating further inductive bias or constraints as its action space can readily be adapted. For instance, the actions for some agents could be explicitly parameterized or coupled with their neighboring agents. In this regard, Closure-RL could also be used to further improve other, domain-specific closure models.
Acknowledgements
S.K. and P.K. acknowledge support by The European High Performance Computing Joint Undertaking (EuroHPC) Grant DCoMEX (956201-H2020-JTI-EuroHPC-2019-1) and support by the Defense Advanced Research Projects Agency (DARPA) through Award HR00112490489.
Appendix A Neural Network Architecture
Figure 5: The IRCNN backbone takes the current global observation and maps it to a feature tensor. This feature tensor is passed into two different convolutional layers that predict the per-discretization point action-distribution-parameters and state-values. In the case of a PDE with three spatial dimensions, the architecture would need to be based on three-dimensional convolutional layers instead.
The optimal policy is expected to compensate for errors introduced by the numerical method and its implementation on a coarse grid. We use the Image Restoration CNN (IRCNN) architecture proposed in [48] as our backbone for the policy- and value-network. A discussion of this choice can be found in Appendix F. In Figure 5 we present an illustration of the architecture and show that the policy- and value network share the same backbone. The policy- and value network only differ by their last convolutional layer, which takes the features extracted by the backbone and maps them either to the action-distribution-parameters $\boldsymbol{\mu}_{A}\in\mathcal{A}$ and $\boldsymbol{\sigma}_{A}\in\mathcal{A}$ or the predicted return value for each agent. The local policies are assumed to be independent, such that we can write the distribution at a specific discretization point as
$\pi_{ij}(\boldsymbol{A}_{ij}|\boldsymbol{O}_{ij})=\mathcal{N}(\boldsymbol{\mu}_{A,ij},\boldsymbol{\sigma}_{A,ij}).$ (11)
Appendix B Adapted PPO Algorithm